What concentration of nitrate ions (in ppm) is present in 7.1 × 10 –7 M calcium nitrate solution? For very dilute aqueous solutions, you can assume the solution's density is 1.0 g/ mL.
Answer #1

ppm = mg of solute / kg of solvent

kg = 1 kg, for 1 liter of solution

mol = MV = (7.1*10^-7)(1) = 7.1*10^-7 mol of Ca(NO3)2

mol of NO3- = 2*7.1*10^-7 mol of Ca(NO3)2 = 14.2 *10^-7 mol

then

mass = mol*MW = ( 14.2 *10^-7 )(62.0049) = 0.000088046 g = 0.088046 mg

ppm = (0.088046 ) mg/kg = 0.088046 ppm